# The impact of stellar bars on quenching star formation: insights from a spatially resolved analysis in the local Universe

### v1

@@ 3.2 Photometry and SED fitting
In order to perform the pixel-by-pixel SED fitting, it is necessary to measure the flux at each photometric band within the apertures. To this purpose, we generate a grid using a customizable process, through which the user can choose the cell size (, in our case) and the radius of the circle, centred on the galaxy centre, within which the grid is built. Then, the fluxes at the various wavelengths are measured in every cell through the photutils v0.6 Python package (Bradley et al., 2019). Those at wavelengths shorter than 10 m are also corrected for Galactic extinction, thanks to the in-built module in CAAPR, based on values in the IRSA Galactic Dust Reddening and Extinction Service (Schlafly & Finkbeiner, 2011). Subsequently, to give an estimate of the flux errors, we use the error maps in the DustPedia archive, which are typically available for the Spitzer bands and for the far-IR photometry. In cases where no error map is accessible (i.e., the SDSS maps), the uncertainty for a given flux is computed as the quadrature sum of the calibration noise and the root mean square (rms) of the maps, similarly to what has been done by the DustPedia collaboration (Clark et al., 2018). The outcome of this step of the data analysis is a photometric catalogue that contains the flux in the available photometric bands from far-UV to far-IR and the corresponding uncertainties, along with the associated astrometric positions of the pixel/cell centres. It is worth mentioning that, for each cell, photometric bands that are characterised by a Signal-to-Noise Ratio (SNR) lower than 3 are rejected. Additionally, if a specific aperture exhibits more than 10 rejected photometric bands due to an excessively low SNR, it is excluded from the subsequent analysis.
@@
As already anticipated, we perform the SED fitting procedure using the magphys code. Through a Bayesian approach, it models the whole UV-to-FIR emission, assuming a balance between the energy emitted at UV/optical/NIR wavelengths that is absorbed by dust and the one re-emitted in the FIR. In order to fit the observed multi-wavelength photometry, magphys relies on the Bruzual & Charlot (2003) stellar population synthesis code, which is used to compute the light produced by stars in galaxies. It predicts the spectral evolution of stellar populations for varying star formation histories, which are described by a continuous model with superimposed random bursts, following Kauffmann et al. (2003). Galaxies are assumed to produce stars according to this law from a certain formation time tform, ranging between and , to the present. The typical values of the star formation time-scale parameter are approximately uniformly distributed over the interval from to , dropping exponentially around . The models also include metallicities ranging from 0.2 to . To characterise the attenuation of starlight by dust, the two-component model proposed by Charlot & Fall (2000) is employed (da Cunha et al., 2008). This accounts for the luminosity absorbed and re-emitted by dust in stellar birth clouds (i.e., giant molecular clouds) and in the diffuse ISM in galaxies. The infrared emission from molecular clouds is characterised by the sum of three components: a component of polycyclic aromatic hydrocarbons (PAHs); a mid-infrared continuum associated with the emission from hot grains at temperatures between 130 and 250 K; and a component of grains in thermal equilibrium with an adjustable temperature in the range of 30 – 60 K. For the diffuse ISM emission, a component of cold grains with temperatures of 15 – 25 K is included, too.
@@
Following Smith & Hayward (2018), the previously measured fluxes (in all the grid apertures) are given as an input to magphys for SED fitting in side cells. The criterion that is adopted to establish if a fit has to be accepted or rejected is based on a cut (Hayward & Smith, 2015; Smith & Hayward, 2018), where such a threshold value is fixed according to the number of photometric bands that are available for the selected objects. In particular, since the galaxies of interest in this work are observed in a number of bands that can vary from 18 to 23, we adopt a conservative threshold value of 25 for the minimum , remaining consistent with the choice made in Enia et al. (2020).
@@ 3.3 Stellar mass and SFR estimates
Finally, the output of this spatially resolved SED fitting process, performed by the magphys code, consists of a wide range of physical properties. Specifically, we are interested in stellar mass () and star formation rate (SFR), which are the two quantities that play a fundamental role in the characterisation of the Main Sequence relation.
Stellar masses are taken from the magphys output, while SFR estimates are obtained by summing the unobscured component (SFRUV) and the one that is obscured by dust absorption and then re-emitted in the IR (SFRIR). Such empirical approach takes into account the contribution of two extreme stellar populations: the UV emission from young stars and the obscured young stellar component that is still hidden in the dusty molecular clouds. The SFRUV contribution is estimated using the relation of Kennicutt (1998):
where is the luminosity at , expressed in erg s-1 Hz-1 and taken from the SED fit. The SFRIR component is also derived from the relation of Kennicutt (1998):
@@
The SFR values are derived from the above empirical relations in order to reduce their dependence on the grid of magphys templates with respect to the ones directly obtained from the SED fit, which are, indeed, discretized and more dependent on degenerate parameters like age, metallicity, and extinction. This choice is also supported by what Enia et al. (2020) found for normal galaxies: the SFR computed as the sum of Equations 1 and 2 and the one given by magphys turned out to be consistent (see Appendix A).
@@ 4 Results
In this Section, we outline the key findings of our analysis, starting with the presentation of the spatially resolved distributions of the stellar mass surface density and of the star formation rate surface density . Figs. 2 and 3 display maps of and in the second and third columns, respectively. All these plots also include: a solid black ellipse of semi-major axis equal to R25, representing the 25-th level isophote of each galaxy, and a dashed black ellipse with the same axis ratio as the 25-th isophote but rescaled in accordance with the bar radius.
Then, we show and analyze the spatially resolved MS relations of all the sources, which are collected in the first column of Figs. 2–3. As it can be seen from these plots, the distribution of points (each one representing the values of and in a certain grid cell) is colour-coded according to the aperture’s normalised distance from the galaxy centre: dnorm. The normalisation of this radial distance is performed with respect to the 25-th level isophote. Specifically, the distance of each cell from the centre is divided by the radial distance of the point lying on the ellipse of the 25-th isophote along the direction identified by that cell. The points that are circled in red correspond to the square apertures lying within the dashed black ellipse, so they are sampling the bar region. In particular, once the bar radius is normalised with the aforementioned procedure (Rbar,norm), we identify these points as the ones characterised by dnorm Rbar,norm.
For each galaxy, the best-fit MS is displayed as a black line, while the MS obtained in the work of Enia et al. (2020) for a sample of unbarred grand design spiral galaxies from the DustPedia archive is reported as a grey line. The latter is given by the following expression:
In the end, the fourth column of Figs. 2–3 collects the spatially resolved maps that are colour-coded as a function of the distance from the best-fit MS of each galaxy. Also in this case, the solid and dashed black ellipses mark the 25-th isophote and the bar region, analogously to what is also represented in and maps.
@@ 4.1 Spatially resolved distributions of and SFR
As already mentioned in Sect. 3.3, the stellar mass values of the single cells are obtained from the magphys output. Then, it is straightforward to get the corresponding surface density by dividing the stellar mass of each grid element by its area. In all the galaxies of our sample, the stellar mass surface density (second column plots of Figs. 2–3) appears to smoothly increase, going from the galaxy outskirts towards the inner regions, as expected for the typical exponential growth of disk galaxies. Additionally, an almost spheroidal structure in the centre is evident in every galaxy and the stellar bar appears to be grafted onto it.
@@
Regarding the distributions of (third column plots of Figs. 2–3), obtained by dividing the sum of SFRUV and SFRIR by each cell area, it can be noticed that the spiral arms are well traced in most of the galaxies, especially in NGC 1566 and NGC 5236. Relatively high values of star formation are typically evident in several peaks along the spiral arms and in some of the very central regions (e.g., the nucleus of NGC 3351). The only exception is NGC 3953, which shows an overall smoother distribution of SFR density. Besides, it is possible to identify a nearly annular and more quiescent structure (lower values of ) that coincides with the region that hosts the bar (the one enclosed by the dashed black ellipse). This is obvious in NGC 4725, but is also detectable in NGC 3351 and NGC 4579. The outermost regions of all the selected galaxies exhibit the lowest star formation activity.
The most striking difference between the spatially resolved distributions of and is that stellar mass clearly traces the presence of the bar, while it is almost never visible in SFR maps.
4.2 Spatially resolved Main Sequence relation
@@
The ultimate goal of this study is to probe the MS relation on sub-kpc scales for our sample of barred galaxies. Accordingly, once the two physical properties of our interest have been explored in their spatial distribution, it is necessary to relate them and analyse the distribution in the – plane (first column plots of Figs. 2–3).
@@
In general, going towards the centre of the galaxies, the stellar mass density increases, as does the density of star formation rate, according to the expected log-linear relation. Nevertheless, the presence of a common decreasing feature arises when looking at the overall trend of the MS in the central part of all these barred galaxies. Indeed, a ‘bending’ of the MS relation is observed in correspondence with the points sampling the region where the bar is located, which are the ones circled in red.
In most cases, the values of clearly tend to increase again at the high end.
@@
For each galaxy, the functional form of the spatially resolved MS relation is obtained by performing a linear fit, through the Orthogonal Distance Regression function implemented in the scipy package of Python, of the values in the – plane, derived for all the cells within that galaxy. In Table 2, we display slopes , zero points (with their corresponding uncertainties) and scatter of the fitted MS relations for all six barred galaxies. The latter is computed as the standard deviation of the residuals along the y-axis with respect to the best-fit MS.
@@
As already mentioned, the best-fit lines are shown as black lines in the first column plots of Figs. 2–3. By comparing the resolved MS relation of each of our barred galaxies with the one obtained for normal spirals (Eq. 3, grey line, Enia et al. (2020)), we observe that the distribution of points typically lies below the relation found for unbarred galaxies, consistent with a systematically lower star formation rate at fixed stellar mass. This is valid for NGC 3351, NGC 3953, NGC 4579 and NGC 4725, while NGC 1566 and NGC 5236 show higher values of , as is expected from the presence of highly star-forming spiral arms and nuclear regions. Furthermore, the common trend of MS bending in the bar-hosting regions appears to anti-correlate with respect to the best-fit relations (black lines), although this effect is more striking in some galaxies than in others.
@@
While the properties discussed so far capture the overall trend of the spatially resolved MS relation of barred galaxies, there are still several galaxy-to-galaxy variations that need to be highlighted. The best-fit MS of NGC 1566 is almost overlapping with the MS of normal spirals and they are consistent within the uncertainties. In NGC 3351, NGC 3953 and NGC 4725, we note that the slopes of the fitted MS relations are similar to those of unbarred spirals, while the zero points have smaller values. On the other hand, both NGC 4579 and NGC 5236 show a steeper slope of the fitted MS than the one found for unbarred galaxies, meaning that the increase of SFR towards the galaxy centre is more rapid. However, while NGC 4579 presents values of that are globally lower than the typical trend of normal galaxies, NGC 5236 stands out as a highly star-forming source.
@@
Focusing on the inner regions of the galaxies, which are the ones of interest, in order to better investigate the impact of the bar structure, we can point out that NGC 4579, NGC 4725 and NGC 5236 (Fig. 3) are the ones with the most evident feature of anti-correlation. Indeed, if the points associated with region that hosts the stellar bar turn out to be arranged in a direction approximately orthogonal to the fitted MS, then the decreases with increasing mass density. Moreover, going towards the very centre of these three galaxies, it is remarkable that the SFR density gets back to having an increasing trend as the stellar mass density grows, but these very central points still appear to lie below the best-fit MS of each galaxy. This implies that the nuclear regions are less star-forming than what would be predicted from based on their own MS. The other three galaxies in our sample, namely NGC 1566, NGC 3351 and NGC 3953 (Fig. 2), exhibit a smoother decline of in the bar region. In particular, the MS of NGC 1566 is also more scattered than the others, reflecting the wider range of values that mostly characterise the spiral arms winding. Besides that, unlike all the other galaxies, in NGC 3351 we can notice that the closest points to the galaxy centre lie above the best-fit relation, thus hinting at a more active star-forming region than what would be foreseen by the MS, given their stellar mass density.
@@ 4.3 Distance from MS
Looking at the panels in the last column of Figs. 2–3, one can get an idea of what is happening in each of these galaxies: the red cells correspond to regions that are less star-forming than what the fitted MS would predict, while the blue ones are associated with particularly active, starbursting regions. For what has been said so far, the central parts (including both the nuclear and the bar hosting regions) of most of the selected objects appear to be red, with suppressed or lower levels of star formation activity compared to the global galaxy MS relation, or even quenched. On the other hand, their spiral arms show a tendency to be active in star formation. On the contrary, NGC 3351 exhibits a star-forming nucleus with a red (quiescent) annular region around it.

### v2

@@ 3.2 Photometry and SED fitting
In order to perform the pixel-by-pixel SED fitting, it is necessary to measure the flux at each photometric band within the apertures. To this purpose, we generated a grid using a customisable process, through which the user can choose the cell size (, in our case) and the radius of the circle centred on the galaxy centre within which the grid is built. The fluxes at the various wavelengths are then measured in every cell using the photutils v0.6 Python package (Bradley et al., 2019). Those at wavelengths shorter than 10 m are also corrected for Galactic extinction based on values in the IRSA Galactic Dust Reddening and Extinction Service thanks to the in-built module in CAAPR (Schlafly & Finkbeiner, 2011). Subsequently, to give an estimate of the flux errors, we used the error maps in the DustPedia archive, which are typically available for the Spitzer bands and for the FIR photometry. In cases where no error map is accessible (i.e. the SDSS maps), the uncertainty for a given flux is computed as the quadrature sum of the calibration noise and the root mean square (rms) of the maps, which is similar to the approach of the DustPedia collaboration (Clark et al., 2018). The outcome of this step of the data analysis is a photometric catalogue that contains the flux in the available photometric bands from FUV to FIR —with the corresponding uncertainties— along with the associated astrometric positions of the cell centres. It is worth mentioning that, for each cell, photometric bands that are characterised by a signal-to-noise ratio (S/N) of lower than 3 are rejected. Additionally, if a specific aperture exhibits more than ten rejected photometric bands due to an excessively low S/N, it is excluded from the subsequent analysis.
@@
As mentioned above, we performed the SED fitting procedure using the magphys code. Through a Bayesian approach, it models the whole UV-to-FIR emission, assuming a balance between the energy emitted at UV, optical, and NIR wavelengths that is absorbed by dust and the emission re-emitted in the FIR. In order to fit the observed multi-wavelength photometry, magphys relies on the Bruzual & Charlot (2003) stellar population synthesis code, which is used to compute the light produced by stars in galaxies. This code predicts the spectral evolution of stellar populations for varying star formation histories, which are described by a continuous model with superimposed random bursts, following Kauffmann et al. (2003). Galaxies are assumed to produce stars according to this law from a certain formation time tform, ranging between and , to the present. The typical values of the star formation timescale parameter are approximately uniformly distributed over the interval from to , dropping exponentially around . The models also include metallicities ranging from 0.2 to . To characterise the attenuation of starlight by dust, the two-component model proposed by Charlot & Fall (2000) is employed (da Cunha et al., 2008). This accounts for the luminosity absorbed and re-emitted by dust in stellar birth clouds (i.e. giant molecular clouds) and in the diffuse ISM in galaxies. The infrared emission from molecular clouds is characterised by the sum of three components: a component of polycyclic aromatic hydrocarbons (PAHs); a mid-infrared continuum associated with the emission from hot grains at temperatures of between 130 and 250 K; and a component of grains in thermal equilibrium with an adjustable temperature in the range of 30 – 60 K. For the diffuse ISM emission, a component of cold grains with temperatures of 15 – 25 K is also included.
@@
Following Smith & Hayward (2018), the previously measured fluxes (in all the grid apertures) are given as an input to magphys for SED fitting in cells of per side. The criterion that is adopted to establish whether or not a fit has to be accepted or rejected is based on a cut (Hayward & Smith, 2015; Smith & Hayward, 2018), where such a threshold value is fixed according to the number of photometric bands that are available for the selected objects. In particular, as the galaxies of interest in this work are observed in a number of bands that can vary from 18 to 23, we adopt a conservative threshold value of 25 for the minimum , remaining consistent with the choice made in Enia et al. (2020).
@@ 3.3 Stellar mass and SFR estimates
Finally, the output of this spatially resolved SED fitting process, performed by the magphys code, consists of a wide range of physical properties. Specifically, we are interested in stellar mass () and SFR, which are the two quantities that play a fundamental role in the characterisation of the MS relation.
Stellar masses are taken from the magphys output, while SFR estimates are obtained by summing the unobscured component (SFRUV) and the component that is obscured by dust absorption and then re-emitted in the IR (SFRIR). This empirical approach takes into account the contribution of two extreme stellar populations: the UV emission from young stars and the obscured young stellar component that is still hidden in the dusty molecular clouds. The SFRUV contribution is estimated using the relation of Kennicutt (1998):
where is the luminosity at expressed in erg s-1 Hz-1 and taken from the SED fit. The SFRIR component is also derived from the relation of Kennicutt (1998):
@@
The SFR values are derived from the above empirical relations in order to reduce their dependence on the grid of magphys templates with respect to the ones directly obtained from the SED fit, which are, indeed, discretised and more dependent on degenerate parameters, such as age, metallicity, and extinction. This choice is also supported by the findings of Enia et al. (2020) regarding normal galaxies: the SFR computed as the sum of Equations 1 and 2 is consistent with the SFR given by magphys (see Appendix A).
@@ 4 Results
In this section, we outline the key findings of our analysis, starting with the presentation of the spatially resolved distributions of the stellar mass surface density and of the SFR surface density . The second and third columns of Figs. 2 and 3 display maps of and , respectively. All these plots also include: a solid black ellipse of semi-major axis equal to R25, representing the 25th level isophote of each galaxy, and a dashed black ellipse with the same axis ratio as the 25th isophote but rescaled in accordance with the bar radius.
We then show the spatially resolved MS relations of all the sources, which are collected in the first column of Figs. 2–3, as well as our analysis of them. As it can be seen from these plots, the distribution of points (each one representing the values of and in a certain grid cell) is colour coded according to the normalised distance of the aperture from the galaxy centre: dnorm. The normalisation of this radial distance is performed with respect to the 25th level isophote. Specifically, the distance of each cell from the centre is divided by the radial distance of the point lying on the ellipse of the 25th isophote along the direction identified by that cell. The points that are circled in red correspond to the square apertures lying within the dashed black ellipse, meaning that they are sampling the bar region. In particular, once the bar radius is normalised with the aforementioned procedure (Rbar,norm), we identify these points as the ones characterised by dnorm Rbar,norm.
For each galaxy, the best-fit MS is displayed as a black line, while the MS obtained in the work of Enia et al. (2020) for a sample of unbarred grand-design spiral galaxies from the DustPedia archive is reported as a grey line. The latter is given by the following expression:
In the end, the fourth column of Figs. 2–3 collects the spatially resolved maps that are colour coded as a function of the distance from the best-fit MS of each galaxy. Also in this case, the solid and dashed black ellipses mark the 25th isophote and the bar region, analogously to what is also represented in and maps.
@@ 4.1 Spatially resolved distributions of and SFR
As already mentioned in Sect. 3.3, the stellar mass values of the single cells are obtained from the magphys output. Then, it is straightforward to get the corresponding surface density by dividing the stellar mass of each grid element by its area. In all the galaxies of our sample, the stellar mass surface density (second column plots of Figs. 2–3) appears to smoothly increase, going from the galaxy outskirts towards the inner regions, as expected for the typical exponential growth of disc galaxies. Additionally, an almost spheroidal structure in the centre is evident in every galaxy and the stellar bar appears to be grafted onto it.
@@
Regarding the distributions of (third column plots of Figs. 2–3) obtained by dividing the sum of SFRUV and SFRIR by each cell area, we note that the spiral arms are well traced in most of the galaxies, and are particularly well traced in NGC 1566 and NGC 5236. Relatively high values of star formation are typically evident in several peaks along the spiral arms and in some of the very central regions (e.g. the nucleus of NGC 3351). The only exception is NGC 3953, which shows an overall smoother distribution of SFR density. In addition, it is possible to identify a nearly annular and more quiescent structure (lower values of ) that coincides with the region that hosts the bar (the one enclosed by the dashed black ellipse). This is obvious in NGC 4725, but is also detectable in NGC 3351 and NGC 4579. The outermost regions of all the selected galaxies exhibit the lowest star formation activity. The most striking difference between the spatially resolved distributions of and is that stellar mass clearly traces the presence of the bar, while it is almost never visible in SFR maps.
4.2 Spatially resolved MS relation
@@
The ultimate goal of this study is to probe the MS relation on subkpc scales for our sample of barred galaxies. Accordingly, after exploring the spatial distribution of the two physical properties of interest here, it is necessary to relate them and analyse the distribution in the – plane (first column plots of Figs. 2–3).
@@
In general, the stellar mass density increases towards the centre of the galaxies, as does the density of the SFR, in agreement with the expected log-linear relation. Nevertheless, the presence of a common decreasing feature arises when looking at the overall trend of the MS in the central part of all these barred galaxies. Indeed, a ‘bending’ of the MS relation is observed in correspondence with the points sampling the region where the bar is located, which are the ones circled in red. In most cases, the values of clearly tend to increase again at the high end.
@@
For each galaxy, the functional form of the spatially resolved MS relation is obtained by performing a linear fit —through the orthogonal distance regression function implemented in the scipy package of Python— of the values in the – plane derived for all the cells within that galaxy. In Table 2, we display slopes , zero points (with their corresponding uncertainties), and scatter of the fitted MS relations for all six barred galaxies. The latter is computed as the standard deviation of the residuals along the y-axis with respect to the best-fit MS.
@@
As already mentioned, the best-fit lines are shown as black lines in the first column plots of Figs. 2–3. By comparing the resolved MS relation of each of our barred galaxies with the one obtained for normal spirals (Eq. 3, grey line; Enia et al. (2020)), we observe that the distribution of points typically lies below the relation found for unbarred galaxies, which is consistent with a systematically lower SFR at fixed stellar mass. This is valid for NGC 3351, NGC 3953, NGC 4579, and NGC 4725, while NGC 1566 and NGC 5236 show higher values of , as expected from the presence of highly star-forming spiral arms and nuclear regions. Furthermore, the common trend of MS bending in the bar-hosting regions appears to be anticorrelated with the best-fit relations (black lines), although this effect is more striking in some galaxies than in others.
@@
While the properties discussed so far capture the overall trend of the spatially resolved MS relation of barred galaxies, there are still several galaxy-to-galaxy variations that need to be highlighted. The best-fit MS of NGC 1566 is almost overlapping with the MS of normal spirals and they are consistent within the uncertainties. In NGC 3351, NGC 3953, and NGC 4725, we note that the slopes of the fitted MS relations are similar to those of unbarred spirals, while the zero points have smaller values. On the other hand, the fitted MS of both NGC 4579 and NGC 5236 shows a steeper slope than that found for unbarred galaxies, meaning that the increase in SFR towards the galaxy centre is more rapid. However, while NGC 4579 presents values of that are globally lower than the typical trend of normal galaxies, NGC 5236 stands out as a highly star-forming source.
@@
Focusing on the inner regions of the galaxies in order to better investigate the impact of the bar structure, we can point out that NGC 4579, NGC 4725, and NGC 5236 (Fig. 3) are the galaxies with the most evident feature of anti-correlation. Indeed, if the points associated with the region that hosts the stellar bar turn out to be arranged in a direction approximately orthogonal to the fitted MS, then the decreases with increasing mass density. Moreover, going towards the very centre of these three galaxies, it is remarkable that the SFR density reverts to having an increasing trend as the stellar mass density grows; however, these very central points still appear to lie below the best-fit MS of each galaxy. This implies that the nuclear regions are less star-forming than what would be predicted from based on their own MS. The other three galaxies in our sample, namely NGC 1566, NGC 3351, and NGC 3953 (Fig. 2), exhibit a smoother decline in in the bar region. In particular, the MS of NGC 1566 is also more scattered than the others, reflecting the wider range of values that mostly characterise the winding spiral arms. In addition, as opposed to the other galaxies, in NGC 3351 we note that the closest points to the galaxy centre lie above the best-fit relation, which suggests a more active star-forming region than what would be foreseen by the MS, given their stellar mass density.
@@ 4.3 Distance from MS
The panels in the last column of Figs. 2–3 provide some information about what is happening in each of these galaxies: the red cells correspond to regions that are more weakly star forming than what the fitted MS would predict, while the blue ones are associated with particularly active, starbursting regions. The central parts (including both the nuclear and the bar hosting regions) of most of the selected objects appear to be red, with suppressed or lower levels of star-formation activity compared to the global galaxy MS relation, or perhaps even quenched star formation. On the other hand, their spiral arms show a tendency to be active in star formation. On the contrary, NGC 3351 exhibits a star-forming nucleus with a red (quiescent) annular region around it.